OpenStudy (winsteria):

A recipe calls for 4 cups of flour. Lauren wants to make 20% more servings than the recipe makes. Which expression represents how many cups of flour Lauren should use? A) 1/5 ∗ 4 B) 4/5 ∗ 4 C) 6/5 ∗ 4 D) 9/5 ∗ 4

OpenStudy (winsteria):

@KamiBug @19199

OpenStudy (aripotta):

so 20% more means 120% right? so if 20% = 1/5 then 120% = 6/5

OpenStudy (winsteria):

So C?

OpenStudy (aripotta):

yep

OpenStudy (winsteria):

Thank you

OpenStudy (aripotta):

no problemo :)